OpenStudy (anonymous):

LS=RS sec^2x - sec^2y = tan^2x-tan^2y

OpenStudy (anonymous):

\[\cos^2(x)+\sin^2(x)=1 \implies 1 + \tan^2(x)=\sec^2(x)\] \[\implies \boxed{\sec^2(x)-\sec^2(y)}=1+\tan^2(x)-1-\tan^2(y)=\boxed{\tan^2(x)-\tan^2(y)}\]

OpenStudy (anonymous):

I have no idea what you did there

OpenStudy (anonymous):

Used the pythagorean identity. Divide by cos^2(x) to rewrite it as the right side of the first equation. Then plug in what you have and cancel the 1s.

OpenStudy (anonymous):

What is the pythagorean identity?

OpenStudy (anonymous):

\[\boxed{\cos^2(x) + \sin^2(x)=1}\]

OpenStudy (anonymous):

How does that apply within sec^2x - sec^2y = tan^2x-tan^2y ?

OpenStudy (anonymous):

What is secant in terms of cosine?

OpenStudy (anonymous):

secx = 1/cosx?

OpenStudy (anonymous):

Exactly. So divide both sides of the pythagorean identity by cos^2(x): \[\frac{\cos^2(x)+\sin^2(x)}{\cos^2(x)} = \frac{1}{\cos^2(x)} \implies 1 + \frac{\sin^2(x)}{\cos^2(x)} = \sec^2(x)\] But what is sin(x)/cos(x)?

OpenStudy (anonymous):

sinx/cosx = tan x?

OpenStudy (anonymous):

Yes. So you have: \[\sin^2(x)+\cos^2(x) = 1 \implies 1+\tan^2(x) = \sec^2(x)\] And you have two sec^2(x) quantities. So replace those with this identity: \[(1+\tan^2(x))-(1+\tan^2(y))=\tan^2(x)-\tan^2(y)\] which is exactly what you wanted to show.

OpenStudy (anonymous):

so sec^2x would = 1+tan^2x?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So you would sub in 1 + tan ^2x/y into both sec^2x/y? what happens to the 1s?

OpenStudy (anonymous):

1-1 = 0

OpenStudy (anonymous):

Don't forget to distribute the negative. \[ - (a+b) = -a-b\]

OpenStudy (anonymous):

ah okay sorry brakets confused me

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh thank you so much :)